shiela has 26 coins in her toy bank. They are all nickels and dimes, and their total value is $2.45. How many dimes and nickels dose she have?
Answer by stanbon(75887)

shiela has 26 coins in her toy bank. They are all nickels and dimes, and their total value is $2.45. How many dimes and nickels dose she have?
-------
Equations:
n + d = 26
5n+10d = 245
--------------------
Modify:
5n + 5d = 5*26
5n +10d = 245
--------------------
Subtract and solve for "d":
5d = 115
d = 23 (# of dimes)
--------
Solve for "n":
n + d = 26
n + 23 = 26
n = 3 (# of nickels)
=================
Cheers,
Stan H.